# Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine-Learning Molecular Dynamics

### v2

@@ IV Conclusions
We have demonstrated that, when applied to the prediction of the IR spectra of PAHs, MLMD provides predictive accuracy comparable to that of traditional quantum chemical methods but at a fraction of the computational cost. With a scaling law of approximately $N_{\text{C}}^{0.7}$, MLMD is particularly suitable for large PAHs, which are key contributors to AIBs. In contrast, traditional quantum chemical anharmonic methods scale much more steeply, approximately $N_{\text{C}}^{4.1}$. Furthermore, we have shown that MLMD is an ideal method for studying temperature effects in molecular IR spectra, which might be critical for understanding astronomical observations. By calculating the anharmonic spectra of 1 704 neutral PAHs from PAHdb at 50, 300, and 600 K, with sizes up to 216 carbon atoms, we highlight that the efficiency of MLMD enables the creation of extensive molecular anharmonic spectral datasets. These datasets could be of instrumental importance for the advancement of AI-assisted astronomical analyses of observational IR spectra in the future.
@@
Despite the impressive efficiency of MLMD, several limitations persist. First, the current model was trained exclusively on neutral PAHs with natural chemical elements and does not account for charged molecules or isotopologues, both of which may play a significant role in astronomical contexts. This limitation stems from the scarcity of data for the IR spectra, as well as challenges in incorporating charge state and isotope information into molecular descriptors. Second, because of the data-driven nature of the approach, the model may exhibit increased uncertainty when predicting spectra for molecules that deviate substantially from those in the training dataset. We are actively working to address these challenges, with the goal of further enhancing MLMD’s capabilities for IR spectrum computation.
Data Availability
@@
The source code and the MLMD model to calculate the anharmonic spectra of PAHs using MLMD are freely available as the supplementary information of this article. The model is trained and ready for use without the need for a further ML procedure. It will be continuously updated to improve predictive performance. The supplementary file also includes the spectral data for the 1 704 theoretically-calculated and 49 experimentally-tested PAHs.
Appendix I
The structures and chemical formulas of the 49 PAHs, whose experimentally measured IR spectra are compared with the theoretical spectra in this study, are presented in Figure 8. The spectra and the corresponding relaxed structures are obtained from version 3.2 of PAHdb (Mattioda et al., 2020).
Figure 8: Structures and chemical formulas of the 49 experimentally measured PAHs.

### v3

@@ 4 Conclusions
We have demonstrated that, when applied to the prediction of the IR spectra of PAHs, MLMD provides predictive accuracy comparable to that of traditional quantum chemical methods but at a fraction of the computational cost. With a scaling law of approximately $N_{\text{C}}^{0.7}$, MLMD is particularly suitable for large PAHs, which are key contributors to AIBs. In contrast, traditional quantum chemical anharmonic methods scale much more steeply, approximately $N_{\text{C}}^{4.1}$. Furthermore, we have shown that MLMD is an ideal method for studying temperature effects in molecular IR spectra, which might be critical for understanding astronomical observations. By calculating the anharmonic spectra of 1,704 neutral PAHs from PAHdb at 50, 300, and 600 K, with molecular sizes containing up to 216 carbon atoms, we highlight that the efficiency of MLMD enables the creation of extensive molecular anharmonic spectral datasets. These datasets could be of instrumental importance for the advancement of AI-assisted astronomical analyses of observational IR spectra in the future.
@@
Despite the impressive efficiency of MLMD, several limitations persist. First, the current model was trained exclusively on neutral PAHs with natural chemical elements and does not account for charged molecules or isotopologues, both of which may play a significant role in astronomical contexts. This limitation stems from the scarcity of data for the IR spectra, as well as challenges in incorporating charge state and isotope information into molecular descriptors. Second, because of the data-driven nature of the approach, the model may exhibit increased uncertainty when predicting spectra for molecules that deviate substantially from those in the training dataset. While MLMD captures a broad spectrum of anharmonic effects through classical sampling of the potential energy surface, it does not account for quantum-specific phenomena such as Fermi resonance, which arise from interactions between quantized vibrational states and require a fully quantum mechanical treatment. Third, the training data of the current model were mainly sourced from molecular conformations at 300 K, which are in principle not sufficient for reliably capturing IR spectral features at very high temperatures (e.g., >600 K). We are actively working to address these challenges, with the goal of further enhancing MLMD’s capabilities for IR spectrum computation.
Supplementary Materials
@@
The source code and the MLMD model to calculate the anharmonic spectra of PAHs using MLMD are freely available at DOI: 10.5281/zenodo.14998197. The model is ready for use without the need for further ML training. It will be continuously updated to improve predictive performance. This repository also includes the spectral data for the 1 704 theoretically-calculated and 49 experimentally-tested PAHs. The Supplementary Information is also included in this repository.